# Effective and Stealthy One-Shot Jailbreaks on Deployed Mobile Vision–Language Agents

### v1

@@ Abstract
Large vision-language models (LVLMs) enable autonomous mobile agents to operate smartphone user interfaces, yet vulnerabilities to UI-level attacks remain critically understudied. Existing research often depends on conspicuous UI overlays, elevated permissions, or impractical threat models, limiting stealth and real-world applicability. In this paper, we present a practical and stealthy one-shot jailbreak attack that leverages in-app prompt injections: malicious applications embed short prompts in UI text that remain inert during human interaction but are revealed when an agent drives the UI via ADB (Android Debug Bridge). Our framework comprises three crucial components: (1) low-privilege perception-chain targeting, which injects payloads into malicious apps as the agent’s visual inputs; (2) stealthy user-invisible activation, a touch-based trigger that discriminates agent from human touches using physical touch attributes and exposes the payload only during agent operation; and (3) one-shot prompt efficacy, a heuristic-guided, character-level iterative-deepening search algorithm (HG-IDA*) that performs one-shot, keyword-level detoxification to evade on-device safety filters. We evaluate across multiple LVLM backends, including closed-source services and representative open-source models within three Android applications, and we observe high planning and execution hijack rates in single-shot scenarios (e.g., GPT-4o: 82.5% planning / 75.0% execution). These findings expose a fundamental security vulnerability in current mobile agents with immediate implications for autonomous smartphone operation.
Warning: This paper contains unsafe agent behaviors.
@@ 1 Introduction
Large vision–language models (LVLMs) have enabled mobile agents that interpret natural-language instructions and autonomously carry out complex tasks on smartphones. These agents perform user-facing actions such as social-media interactions, routine note taking, and smart-home control. These capabilities are swiftly evolving from research prototypes to commercial applications (Wang et al., 2024c; Zhang et al., 2025a; Wu et al., 2024a; Yang et al., 2025; Hu et al., 2025). Emerging agent frameworks (Jiang et al., 2025; Wang et al., 2025b) integrate multiple specialized sub-agents to manage reasoning-intensive, long-horizon workflows that reflect real-world complexity. However, because these agents operate on sensitive on-device data and can autonomously initiate real-world actions, successful compromises can lead to severe consequences, ranging from privacy violations and financial losses to safety risks. Moreover, the integration of visual UI perception with language-driven planning introduces unprecedented interaction-level vulnerabilities that go beyond traditional smartphone security paradigms.
Recent works have begun to exploit these new attack surfaces, but they exhibit significant practical limitations. First, while explicit GUI manipulations (such as transparent overlays or deceptive pop-ups) can mislead agent perception or intercept simulated taps (Wu et al., 2025), these techniques remain visually conspicuous, making them impractical in real-world human-occupied environments. Second, although tampering with interface elements through poisoned icons or cloned UI assets can corrupt an agent’s input pipeline, they presuppose that adversaries possess the ability to place or control UI resources (Yang et al., 2024). This significant privilege limits the feasibility for attackers with restricted system access. Third, OS-level injection channels can deliver adversarial prompts from within the mobile stack (Chen et al., 2025), but these approaches typically require elevated permissions and demonstrate limited resilience against on-device LVLM safety filters. Moreover, many state-of-the-art jailbreak and adversarial techniques rely on multi-turn interactions or iterative optimization (Ha et al., 2025), which are impractical for single-interaction, length-constrained contexts typical of mobile agents. Therefore, current approaches have not simultaneously achieved imperceptibility to human users, deployability without elevated privileges, and single-attempt effectiveness against realistic on-device filtering mechanisms.
@@
To address these shortcomings, we aim to develop a low-privilege, stealthy jailbreak framework that crafts one-shot prompt injections against LVLM mobile agents. This undertaking presents three fundamental technical challenges. First, since mobile agents rely on visual UI snapshots for decision-making, an effective attack must manipulate the agent’s perceived interface within standard permission boundaries. It alters what the agent observes without relying on elevated OS permissions such as overlays, notification listeners, accessibility services, or root access. Such privileges are typically impractical to obtain, readily detectable, and difficult to deploy at scale. Second, malicious content should remain imperceptible during normal human interaction yet become exposed precisely when the agent operates. This requires a covert activation mechanism that discriminates between agent-driven and human input without generating persistent or conspicuous UI artifacts. Third, mobile-agent interactions impose additional constraints through single-turn exchanges with strict screen-space limits and on-device safety classifiers. These constraints require that injections be length-bounded and robust to opaque moderation mechanisms within a single inference. Addressing these three constraints jointly—visual plausibility, selective activation, and one-shot robustness—defines the design space for practical, real-world jailbreaks on mobile agents.
In this paper, we propose a unified attack framework composed of three synergistic components. Low-Privilege Perception-Chain Targeting: to avoid reliance on elevated system permissions, we embed jailbreak prompts entirely within the malicious app’s own UI as notification-like elements rendered during agent interaction. These in-app banners are designed to mimic benign UI affordances so that they are captured by the agent’s screenshot-based perception pipeline while requiring no extra OS privileges or overlays. Stealthy User-Invisible Activation: to preserve stealth, we exploit measurable differences between automated agent inputs (e.g., ADB-driven taps) and human touches. We develop a lightweight trigger detector that monitors input event features (such as touch size and pressure) and conditions prompt display on signals characteristic of automated control, thereby exposing the injected content only in agent-driven execution contexts. One-Shot Prompt Efficacy: to operate within single-turn, length-constrained interactions and to evade opaque on-device filters, we design a character-level detoxification pipeline that produces minimal semantic-preserving perturbations of harmful tokens. Concretely, we introduce HG-IDA*, a heuristic-guided iterative-deepening A* search that selects targeted character edits to key tokens and inducing prefixes, optimizing a tradeoff between safety-score improvement and semantic similarity; the resulting one-shot prompts retain their intent for the agent while reducing detection by built-in LVLM classifiers. Together, these components form a complete pipeline that embeds malicious prompts in a low-privilege manner, reveals them selectively under automated operation, and preserves attack intent while increasing the likelihood of bypassing on-device safety checks in a single inference. Figure 1 illustrates a representative privacy-leakage case.
@@
To evaluate our framework, we develop three representative Android applications and release a dataset of jailbreak‑prompt injections, including explicit harmful prompts and seemingly benign prompts that nonetheless induce malicious behavior in agents, which covers privacy leakage, safety harms, potential financial loss, and illicit IoT control across real app scenarios (social, personal notes, smart-home). Using diverse injection instances, we evaluate Mobile-Agent-E with multiple LVLM backends, including state-of-the-art closed-source models (e.g., GPT-4o (Hurst et al., 2024), Gemini-2.0-pro (DeepMind, 2024)) and advanced open-source models (e.g., Deepseek-VL2 (Wu et al., 2024c), Llava-OneVision (Li et al., 2025)). Our Specificity‑Aware Trigger Detector achieved 100% accuracy in distinguishing agent‑driven ADB interactions from human touch events as shown in Appendix A. In terms of attack efficacy, we observed high attack success rates on both closed- and open-source LVLMs (e.g., 82.5% for GPT-4o and 87.5% for Deepseek-VL2) through comprehensive experiments. Moreover, high-capability closed-source models were more likely to convert compromised plans into executed harmful actions due to stronger reasoning-to-action consistency and superior instruction-following. These results underscore the practicality and robustness of stealthy, one‑shot jailbreak prompt injections against real‑world mobile LVLM agents.
@@ 2 Related Work
Mobile agents. The emergence of mobile LLM agents has enabled autonomous task execution on smartphones via visual‑linguistic reasoning. AppAgent (Zhang et al., 2025b) introduced a multimodal framework that controls Android apps through LLM‑generated action plans based on GUI screenshots. Mobile‑Agent (Wang et al., 2024b) and its extension Mobile‑Agent‑V (Wang et al., 2025a) further improved robustness by incorporating action correction and multi‑agent collaboration. Furthermore, Mobile‑Agent‑E (Wang et al., 2025b) integrates multiple specialized sub‑agents (separating perception, planning, and execution) to handle reasoning‑intensive, long‑horizon tasks more effectively. This modular design makes Mobile‑Agent‑E particularly well suited for automating complex, real‑world smartphone workflows under diverse UI conditions. Other agents, such as InfiGUIAgent (Liu et al., 2025), ClickAgent (Hoscilowicz et al., 2024), and Mobile‑Agent‑V2 (Wang et al., 2024a), share a similar architecture, combining vision‑language models with system‑level APIs to simulate human interactions on mobile devices.
@@
Security of multimodal mobile agents. Extensive research has exposed agent vulnerabilities in non-mobile settings: web and desktop agents are susceptible to prompt-injection attacks that embed adversarial text into pages or dialogs (e.g., WIPI (Wu et al., 2024b); EIA (Liao et al., 2024)). By contrast, the security of mobile vision–language agents has only recently attracted attention: (Wu et al., 2025) performed a systematic attack‑surface analysis and demonstrate GUI-based hijacks such as transparent overlays and pop‑up dialogs to mislead agent perception. However, these attacks rely on overt UI changes requiring overlay permissions and lack covert triggering strategies.  (Yang et al., 2024) proposed a systematic security matrix and showcased adversarial UI elements, including poisoned icons and manipulated screenshots. While insightful, their threat model assumes full control over UI assets and does not account for agent behavior under realistic execution constraints. (Chen et al., 2025) introduced the Active Environment Injection Attack (AEIA), in which malicious prompts are injected via system notifications to influence agent decisions. While effective in interrupting agent workflows, AEIA depends on privileged access to notification channels and does not demonstrate success in bypassing LLM safety filters. To our knowledge, none of these studies investigate low-privilege, stealthy, and one-shot jailbreaks under practical UI constraints.
@@
Jailbreak attacks. Prior research can be grouped into two complementary strands. On the one hand, single-shot, non-iterative techniques have shown that carefully designed prefixes or contextual role-plays can subvert alignment constraints—for example, the “Do Anything Now” (DAN) family systematically induces models to ignore safety guards (Shen et al., 2024). In white-box settings, optimization-based methods such as GCG (Zou et al., 2023) craft adversarial suffixes via gradient signals; these suffixes can be generated offline and applied in a one-shot, transferable manner. On the other hand, automated jailbreak generators (e.g., AutoDAN (Liu et al., 2023), GPTFuzz (Yu et al., 2023)) depend on multi-step search, large query budgets, or stronger access (white-box gradients or external LLM evaluators), and thus are incompatible with our strict one-shot threat model we adopt. Overall, our jailbreak framework for mobile agents jointly addresses low-privilege operation, stealth, and one-shot effectiveness: (i) influences agents’ visual input via in-app prompt injection without elevated permissions, (ii) activates only under agent-driven interactions, and (iii) aims to bypass on-device safety checks in a single inference.
@@ 3 Methodology
3.1 Threat Model and Assumptions
This paper focuses on tricking the agent into performing the attacker-specified malicious instructions rather than the user’s commands. Therefore, we model attackers with the capability to modify an app’s source code but without system-level privileges (no overlay permissions or notification access, and no root privileges). This threat model reflects realistic scenarios where developers or maintainers could introduce malicious modifications. The target mobile agent (e.g., Mobile‑Agent‑E) operates via ADB‐driven touch events in line with emerging agent frameworks. Our attack embeds a one‑shot jailbreak prompt entirely within the malicious app’s UI and employs a covert trigger mechanism that reveals it only when under agent control. This design enables workflow hijacking across multiple apps without requesting additional permissions. This approach differs fundamentally from previous GUI-overlay attacks that rely on conspicuous UI changes or notification access, and from threat models requiring full control over UI assets. Our framework achieves stealthy in-app prompt injection and bypasses on-device LLM safety mechanisms in a single inference cycle.
3.2 Problem Formalization
Let $\mathcal{S}$ denote the space of UI perception states and $\mathcal{A}$ the agent’s action space, including taps, swipes, text input, etc. A benign agent policy $\pi\colon\mathcal{S}\to\mathcal{A}$ executes a sequence $(a_{1},\dots,a_{T})$ to achieve a user‐specified goal $G_{u}$. In our attack scenario, attackers embed a single‐turn prompt $\delta$ into app’s UI to induce agent to execute the attacker’s goal $G_{a}$. This prompt is never visible under manual interaction but is revealed when the agent drives the UI via ADB. To capture this behavior, let each touch event be
$e_{t}=(x_{t},y_{t},size_{t},pressure_{t}),$
here $e_{t}$ denotes the touch-event vector at time $t$, containing the screen coordinates $(x_{t},y_{t})$, the contact size $size_{t}$, and the touch pressure $pressure_{t}$. Define a binary trigger function
$\chi(e_{t})=\begin{cases}1,&\text{if }e_{t}\text{ is ADB‑driven (agent control)},\\ 0,&\text{otherwise (human control)}.\end{cases}$
We then model the perceived state at step $t$ as
$s^{\prime}_{t}\;=\;s_{t}\;+\;\chi(e_{t-1})\,\delta,$
so that $\delta$ is appended to the UI state only when $\chi(e_{t-1})=1$.
The attack’s objective is to optimize a prompt $\delta^{*}$ (with $\lvert\delta^{*}\rvert\leq L_{\max}$, $S(\delta^{*})\geq\eta$) to maximize the probability that the agent’s action at injection time $t^{*}$ realizes the attacker’s intent, i.e.
$\max_{\delta^{*}:\;|\delta^{*}|\leq L_{\max}}\Pr\!\bigl[\pi(s^{\prime}_{t^{*}})\in\mathcal{A}(G_{a})\ \bigr],$
where $S(\cdot)\in[0,1]$ is the safety score and $\eta\in[0,1]$ is the acceptance threshold, $s^{\prime}_{t^{*}}=s_{t^{*}}+\chi(e_{{t^{*}}-1})\,\delta^{*}$ and $\mathcal{A}(G_{a})$ denotes actions fulfilling the attacker’s intent.
3.3 Perception‑Chain Targeting
Mobile agents largely rely on their captured UI snapshot, so our attack injects the jailbreak prompt as an in‑app notification banner rather than a separate overlay. We implement a notification‑like view at the top of the malicious app’s interface—styled identically to benign toasts. Upon detecting an ADB‑driven tap, the malicious app displays a toast‑style banner at the top of its interface for a preset duration $t^{\prime}$, carrying the attacker’s instruction. Because this banner uses only standard UI APIs within the app, no extra permissions are required, and human users perceive no lasting change while the agent’s next screenshot captures the injected prompt.
3.4 User‑Invisible Activation
Stealthily revealing the jailbreak prompt only during automated agent control is critical to avoid alerting human users. To this end, we detect ADB-driven taps using the trigger function, since these taps typically exhibit near-zero contact size or pressure:
$\chi(e_{t})=\begin{cases}1,&size_{t}\leq\epsilon_{s}\;\lor\;pressure_{t}\leq\epsilon_{p},\\ 0,&\text{otherwise},\end{cases}$
where $e_{t}=(x_{t},y_{t},size_{t},pressure_{t})$ and $\epsilon_{s},\epsilon_{p}$ are small constants (e.g., 0.01, 0.05). We then condition the prompt injection on the previous event by updating the perceived state as
$s^{\prime}_{t}=s_{t}+\chi(e_{t-1})\,\delta,$
so that $\delta$ appears only when $\chi(e_{t-1})=1$. During manual interaction ($\chi=0$), $s^{\prime}_{t}=s_{t}$ and the UI remains unchanged.
3.5 One‑Shot Prompt
In contrast to LLM Jailbreak, LLM-based mobile agents cannot engage in multi‑turn dialogues with injected prompts, and the limited screen real estate imposes a strict character budget $L_{\max}$. Existing jailbreak techniques typically rely on iterative exchanges or unconstrained prompt length, making them impractical on smartphones. To robustly compromise Mobile Agent in a single inference, we must craft a standalone, length‑bounded prompt that both entices the agent to obey the attacker’s command and slips past on‑device safety filters. Therefore, our jailbreak prompt must (1) guide the agent toward the attacker’s intent, (2) evade on‑device safety filters, and (3) respect length limits. We achieve this via a two‑step pipeline: (i) selection and templating of an inducive prompt, and (ii) character‑level detoxification of harmful keywords.

### v2

# Practical and Stealthy Touch-Guided Jailbreak Attacks on Deployed Mobile Vision–Language Agents

@@ Abstract
Large vision-language models (LVLMs) enable autonomous mobile agents to operate smartphone user interfaces, yet vulnerabilities in their perception and interaction remain critically understudied. Existing research often relies on conspicuous overlays, elevated permissions, or unrealistic threat assumptions, limiting stealth and real-world feasibility. In this paper, we introduce a practical and stealthy jailbreak attack framework, which comprises three key components: (i) non-privileged perception compromise, which injects visual payloads into the application interface without requiring elevated system permissions; (ii) agent-attributable activation, which leverages input attribution signals to distinguish agent from human interactions and limits prompt exposure to transient intervals to preserve stealth from end users; and (iii) efficient one-shot jailbreak, a heuristic iterative deepening search algorithm (HG-IDA*) that performs keyword-level detoxification to bypass built-in safety alignment of LVLMs. Moreover, we developed three representative Android applications and curated a prompt-injection dataset for mobile agents. We evaluated our attack across multiple LVLM backends, including closed-source services and representative open-source models, and observed high planning and execution hijack rates (e.g., GPT-4o: 82.5% planning / 75.0% execution), exposing a fundamental security vulnerability in current mobile agents and underscoring critical implications for autonomous smartphone operation.
Practical and Stealthy Touch-Guided Jailbreak Attacks on Deployed Mobile Vision–Language Agents
Renhua Ding1,2, Xiao Yang2, Zhengwei Fang2, Jun Luo2, Kun He1, Jun Zhu2 1 School of Computer Science, Huazhong University of Science and Technology 2 Dept. of Comp. Sci. and Tech., Institute for AI, Tsinghua-Bosch Joint ML Center, THBI Lab, BNRist Center, Tsinghua University renhua@hust.edu.cn, yangxiao19@tsinghua.org.cn
@@ 1 Introduction
Large vision–language models (LVLMs) have enabled mobile agents that interpret natural-language instructions and autonomously carry out complex tasks on smartphones (survey_1; survey_2). These agents perform user-facing actions such as social-media interactions, routine note taking, and smart-home control. These capabilities are swiftly evolving from research prototypes to commercial applications (AppAgentX; yang2025mla; os_agents). Emerging agent frameworks integrate multiple specialized sub-agents to manage reasoning-intensive, long-horizon workflows that reflect real-world complexity (survey_3; mobile_agent_e). However, because these agents operate on sensitive on-device data and can autonomously initiate real-world actions, successful compromises can lead to severe consequences, ranging from privacy violations and financial losses to safety risks. Therefore, understanding and mitigating the security vulnerabilities inherent to such agents is critical for ensuring their safe and trustworthy deployment.
Prior efforts have begun probing security risks in mobile agents, but existing attack strategies face substantial practical constraints. For example, explicit visual manipulations such as transparent overlays can mislead agent perception (1), yet they always remain conspicuous in real-world environments, easily noticed by users; attacks that modify interface elements such as poisoned icons (yang2024) typically assume adversarial control over UI resources, but a privilege rarely available to attackers with limited system access; moreover, OS-level injection channels (chen2025) often require elevated permissions and show limited robustness against built-in safety alignment of LVLM. Taken together, existing works fail to provide a practical attack surface, achieving the combination of stealthiness, low-privilege deployability, and Jailbreak effectiveness required for real-world mobile-agent settings.
Although re demonstrate that small, human-imperceptible image perturbations can reliably hijack agent behavior, they also report that such perturbations transfer poorly to unseen models, especially to commercial closed-source LVLMs. This limitation makes the approach impractical in typical mobile agent settings, which predominantly rely on proprietary LVLM backends. Hence, we ensure stealth not by making the injected prompts visually imperceptible to humans, but by restricting their visibility to a brief interval within the agent’s perception phase during the task cycle to reduce the risk of detection from users. In addition, most vision–language mobile agents reason and plan from the visual screenshot of the current step, so it is particularly important for attackers to achieve a successful jailbreak within a one-shot interaction in this setting (oneshot1).
In this work, we present a practical and stealthy one-shot jailbreak attack framework tailored for LVLM-driven mobile agents, composed of three synergistic components. (i) Non-privileged perception compromise: embeds visual jailbreak payloads directly within the malicious application interface, enabling injection without elevated system permissions. (ii) Agent-attributable activation: leverages input attribution signals to distinguish agent generated interactions from human touches, ensuring that injected prompts remain invisible to end users while appearing during agent perception phase. (iii) Efficient one-shot jailbreak: introduces HG-IDA*, a heuristic iterative-deepening A* search algorithm that performs targeted keyword-level detoxification to evade LVLM safety filters while preserving jailbreak intent. These components form a unified pipeline enabling low-privilege, agent-only, and effective one-shot jailbreaks.
@@
To evaluate our framework, we develop three Android applications and release a dataset of jailbreak‑prompt injections, including explicit harmful prompts and seemingly benign prompts that nonetheless induce malicious behavior in agents, which covers privacy leakage, safety harms, potential financial loss, and illicit IoT control. Using diverse injection instances, we evaluate Mobile-Agent-E with multiple LVLM backends, including state-of-the-art closed-source models (e.g., GPT-4o (gpt-4o) ) and advanced open-source models (e.g., Deepseek-VL2 (deepseek-vl2) ). We observed high attack success rates on both closed- and open-source LVLMs (e.g., 82.5% for GPT-4o and 87.5% for Deepseek-VL2) through comprehensive experiments. Moreover, high-capability closed-source models were more likely to convert compromised plans into executed harmful actions due to stronger reasoning-to-action consistency and superior instruction-following. These results underscore the practicality and robustness of stealthy, one-shot jailbreak prompt injections against real‑world mobile LVLM agents.
@@ 2 Related Work
Mobile agents. The emergence of mobile LLM agents has enabled autonomous task execution on smartphones via visual‑linguistic reasoning. AppAgent (AppAgent) introduced a multimodal framework that controls Android apps through LLM‑generated action plans based on GUI screenshots. Mobile‑Agent (mobile_agent) and its extension Mobile‑Agent‑V (mobile_agent_v) further improved robustness by incorporating action correction and multi‑agent collaboration. Furthermore, Mobile‑Agent‑E (mobile_agent_e) integrates multiple specialized sub‑agents (separating perception, planning, and execution) to handle reasoning‑intensive, long‑horizon tasks more effectively. This modular design makes Mobile‑Agent‑E particularly well suited for automating complex, real‑world smartphone workflows under diverse UI conditions. Other agents, such as InfiGUIAgent (liu2025infiguiagent), ClickAgent (clik_agent), and Mobile‑Agent‑V2 (mobile_agent_v2), share a similar architecture, combining vision‑language models with system‑level APIs to simulate human interactions on mobile devices.
@@
Security of multimodal mobile agents. Extensive research has exposed agent vulnerabilities in non-mobile settings: web and desktop agents are susceptible to prompt-injection attacks that embed adversarial text into pages or dialogs (e.g., WIPI (wu2024wipi); EIA (liao2024eia)). By contrast, the security of mobile vision–language agents has only recently attracted attention: (1) performed a systematic attack‑surface analysis and demonstrate GUI-based hijacks such as transparent overlays and pop‑up dialogs to mislead agent perception. However, these attacks rely on overt UI changes requiring overlay permissions and lack covert triggering strategies.  (yang2024) proposed a systematic security matrix and showcased adversarial UI elements, including poisoned icons and manipulated screenshots. While insightful, their threat model assumes full control over UI assets and does not account for agent behavior under realistic execution constraints. (chen2025) introduced the Active Environment Injection Attack (AEIA), in which malicious prompts are injected via system notifications to influence agent decisions. While effective in interrupting agent workflows, AEIA depends on privileged access to notification channels and does not demonstrate success in bypassing LLM safety filters. To our knowledge, none of these studies investigate low-privilege, stealthy, and one-shot jailbreaks under practical UI constraints.
@@
Jailbreak attacks. Prior research can be grouped into two complementary strands. On the one hand, single-shot, non-iterative techniques have shown that carefully designed prefixes or contextual role-plays can subvert alignment constraints—for example, the “Do Anything Now” (DAN) family systematically induces models to ignore safety guards (shen2023dan). In white-box settings, optimization-based methods such as GCG (zou2023gcg) craft adversarial suffixes via gradient signals; these suffixes can be generated offline and applied in a one-shot, transferable manner. On the other hand, automated jailbreak generators (e.g., AutoDAN (liu2023autodan), GPTFuzz (yu2023gptfuzzer)) depend on multi-step search, large query budgets, or stronger access (white-box gradients or external LLM evaluators), and thus are incompatible with our strict one-shot threat model we adopt. Overall, our jailbreak framework for mobile agents jointly addresses low-privilege operation, stealth, and one-shot effectiveness: (i) influences agents’ visual input via in-app prompt injection without elevated permissions, (ii) activates only under agent-driven interactions, and (iii) aims to bypass on-device safety checks in a single inference.
@@ 3 Methodology
3.1 Threat Model
This paper studies attacks that cause a mobile agent to execute attacker-specified instructions instead of the user’s requests. Accordingly, we adopt a threat model in which attackers can modify an application’s source code but lack system-level privileges. Moreover, the target mobile agent interacts with applications through ADB-driven touch events, as adopted in Emerge mobile-agent frameworks (e.g., Mobile-Agent-E). This threat model reflects more realistic attack vectors, as anyone can convert a benign Android application into an inconspicuous malicious variant by adding only a small amount of lightweight trigger code (see Appendix B) without altering its original architecture. The repackaged app can then be distributed through open-source repositories or third-party platforms. For example, an attacker may release a seemingly more feature-rich note-taking application that behaves entirely normally for human users but exposes malicious in-app prompts only when operated by a mobile agent, leading to potential privacy leakage without requiring elevated system privileges. Figure 1 illustrates a representative privacy leakage case.
3.2 Framework Overview
Our attack aims to inject a compact jailbreak prompt $\delta$ into the in-app interface such that the prompt becomes visible only when the mobile agent interacts with the application, while remaining inert during normal human usage. The injected prompt modifies the agent’s perceived screenshot at a designated step $t^{\star}$, thereby influencing the policy output.
Let $s_{t}$ denote the agent’s visual observation, and let $e_{t}$ denote the interaction event at step $t$. When the activation condition determined by $e_{t-1}$ is satisfied, the application renders the embedded payload $\delta$, producing a modified observation
$s^{\prime}_{t}=\mathrm{Act}(s_{t},\delta,e_{t-1}),$ (1)
where $\mathrm{Act}(\cdot)$ is an in-app activation operator that conditionally integrates $\delta$ into the screenshot. Let $\mathcal{A}_{G_{a}}\subseteq\mathcal{A}$ denote the set of actions aligned with the attacker’s goal $G_{a}$. The attacker seeks a payload $\delta$ that maximizes the probability that the agent, upon observing the modified screenshot $s^{\prime}_{t^{\star}}$, outputs an attacker-desired action:
$\delta^{\star}=\arg\max_{\delta\in\mathcal{D}}\Pr\!\left[\pi(s^{\prime}_{t^{\star}})\in\mathcal{A}_{G_{a}}\;\middle|\;e_{t^{\star}-1}=1\right],$ (2)
where $\mathcal{D}$ denotes the feasible space of in-app textual payloads, and $e_{t-1}$ is a binary variable with $e_{t-1}=1$ denoting an agent-driven interaction. This unified formulation provides the objective for the entire attack pipeline. The following subsections instantiate the activation operator $\mathrm{Act}(\cdot)$ by explaining how the payload is embedded, triggered, and optimized.
3.2.1 Non-Privileged Perceptual Compromise
We embed the jailbreak prompt $\delta$ directly into standard in-application visual elements rather than relying on overlays, accessibility services, or system-level rendering hooks. Because these elements are intrinsic to the application’s rendering hierarchy, adding or modifying them does not require elevated permissions or special capabilities, allowing our method to achieve perceptual compromise without any system privileges. The attacker needs only control over the application itself, which makes the embedding of the hidden payload $\delta$ feasible without departing from ordinary application behavior.
3.2.2 Agent-attributable Activation
As mentioned above, maintaining strict human-imperceptibility of injected prompts is unrealistic in practical mobile-agent settings (re). Hence, we redefine stealth as limiting prompt exposure to a brief period during the mobile agent’s perception phase, so that it is detectable by the agent while remaining largely imperceptible to human users. To realize this activation condition, we distinguish agent-driven interactions using raw touch profiles. Let
$r_{t}=(size_{t},pressure_{t})$ (3)
denote the raw interaction record at step $t$. We map this record to an agent-attributable event indicator via a simple classifier
$e_{t}=f(r_{t})=\begin{cases}1,&size_{t}\leq\epsilon_{s}\;\lor\;pressure_{t}\leq\epsilon_{p},\\[3.0pt] 0,&\text{otherwise},\end{cases}$ (4)
where $\epsilon_{s}$ and $\epsilon_{p}$ are small thresholds characterizing the near-zero contact profiles of ADB-generated taps.
We then apply the activation operator defined in Eq. 1, which conditions prompt injection on the previous event. When $e_{t-1}=0$ (human interaction), the perceived state remains unchanged ($s^{\prime}_{t}=s_{t}$). When $e_{t-1}=1$, the embedded payload $\delta$ is rendered into the screenshot for exactly one step, ensuring that exposure occurs only within a narrowly timed activation window aligned with the agent’s control flow.
3.2.3 Efficient One-shot Jailbreak
Unlike prior LLM jailbreak work that can iteratively refine prompts through multi-turn interaction with the model, mobile agents typically offer the attacker a single opportunity to inject text into the agent’s perception pipeline. Because the attacker must embed the malicious prompt into the application ahead of time, the prompt cannot be updated or adapted once the application is deployed and used by the victim. In addition, the amount of text that can be rendered on a mobile screen is inherently limited. Under these constraints, achieving a successful jailbreak within a single turn and within the $L_{\max}$ budget becomes essential. To this end, we adopt a two-stage pipeline: (1) templating and selecting an inducing prompt that composes a small set of atomic components into a coherent, action-steering template; and (2) character-level detoxification that minimally perturbs tokens to improve the prompt’s safety score $S(\cdot)$ while preserving semantiSc similarity to the original intent.